### v1

@@ 5 Object Detection in Autonomous Vehicles
[248] TransFusion 2022 CVPR 3D CNN Point 70.2 65.5 86.2 56.7 66.3 58.8 86.1 68.3 44.2
@@
[249] VMVF 2022 CVPR 3D CNN Range 66.3 58.3 84.4 51.5 58.3 53.9 85.3 63.0 32.5
@@
[250] CenterPoint 2021 CVPR VoxelNet Point 65.5 58.0 84.6 51.0 60.2 53.2 83.4 53.7 28.7
@@
[251] Pointformer 2021 CVPR Pointformer Point 63.8 53.8 82.3 45.1 48.3 45.4 85.1 55.7 25.8
@@
[252] CVCNET 2020 NeurIPS 3D CNN Voxel 66.6 58.2 82.6 49.5 59.4 51.1 83.0 61.8 38.8
@@
[228] 3DSSD 2020 CVPR SA Point 56.4 42.6 81.2 47.2 61.4 30.5 70.2 36.0 8.6
@@
[253] 3DVID 2020 CVPR PMPNet Pillar 53.1 45.4 79.7 33.6 47.1 43.1 76.5 40.7 19.9
@@
[254] CBGS 2019 CVPR Sparse CNN Voxel 63.3 52.8 81.1 48.5 54.9 42.9 80.1 51.5 22.3
@@
[255] PointPillars 2019 CVPR 2D CNN Pillar 45.3 30.5 68.4 23.0 28.2 23.4 59.7 27.4 1.1
@@
Table 19 presents a comparative analysis of 3D object detection methods using the LiDAR-based NuScenes dataset. From Table 19, we provide a detailed discussion of three selected models: PointPillars [255], recognized as one of the pioneer model in this field; Voxel Mamba [236], which achieved the highest results in terms of NDS (NuScenes Detection Score), mAP, and trailer detection on the NuScenes dataset; and LION [205], obtained the highest results for car, truck, bus, pedestrian, and cyclist detection on NuScenes dataset. PointPillars [255] adopts a pillar-based representation to enable a purely 2D convolutional pipeline for real-time 3D object detection. The approach first discretizes the point cloud into vertical columns (pillars) with infinite extent along the z-axis, then uses a simplified PointNet to extract point-wise features within each pillar. These are aggregated and scattered into a dense pseudo-image in the bird’s-eye-view plane. A 2D CNN backbone, composed of a top-down path and a multi-scale upsampling path, processes this pseudo-image to extract high-level spatial features. Finally, an SSD-style detection head performs classification and oriented 3D bounding box regression directly in BEV space, allowing end-to-end training without any 3D convolutions.
@@
Voxel Mamba [236] introduces a group-free state-space backbone for 3D object detection, exploiting the linear-time sequence modeling of State Space Models (SSMs) to bypass the computational bottlenecks of Transformer-based voxel encoders. Instead of partitioning the voxel grid into smaller local groups, it serializes all non-empty voxels into a single sequence using a Hilbert Input Layer, which maps 3D voxel coordinates into a 1D order while preserving spatial locality. The backbone employs a Dual-scale SSM Block (DSB), wherein a forward branch processes the high-resolution sequence. In contrast, a backward branch operates on a down-sampled BEV representation, extending the receptive field without quadratic complexity. An Implicit Window Partition (IWP) injects relative position embeddings for intra- and inter-window offsets, enabling the model to recover local context cues without explicit grouping. LION [205] replaces transformer-based voxel backbones with linear recurrent neural operators to model long-range dependencies across large voxel regions while avoiding quadratic complexity. Point clouds are voxelized and organized into fixed-size 3D windows, within which LION Blocks process voxel features hierarchically. Each block contains a pair of linear group RNN operators, applied along orthogonal spatial axes, to propagate information efficiently across extended spatial extents. Before this recurrent mixing, a 3D spatial descriptor based on submanifold convolutions enriches local geometric context to offset the positional information loss during voxel flattening. Within its hierarchy, the model reduces voxel resolution at selected stages, representing features at multiple scale, and giving them higher-level layers access.
@@
Table 20 presents a comparative analysis of 3D object detection methods using LiDAR-based data from the VOD dataset. From Table 20, we provide a detailed discussion of three selected models: ELMAR [256] and MoRAL [257], which achieved the highest results in terms of $AP\%$ for entire area and driving corridor on VOD dataset; and L4DAR [258], recognized as one of the notable model in the context of 3D object detection.
@@
[256] ELMAR 2025 ArXiv DS+SA Point 76.41 69.34 78.91 74.89 91.74 81.35 93.01 88.70
@@
[257] MoRAL 2025 ArXiv Sparse CNN Pillar 71.23 69.67 79.01 73.30 90.91 78.90 96.25 88.68
@@
[259] MutualForce 2025 ICASSP CNN+RPN Pillar 71.67 66.26 77.35 71.76 92.31 76.79 89.97 86.36
@@
[258] L4DR 2024 AAAI PointNet++ Pillar 69.10 66.20 82.80 72.70 90.80 76.10 95.50 87.47
@@
[260] CM-FA 2024 ICRA SA Point 71.39 68.54 76.60 72.18 90.91 80.78 87.80 86.50
@@
[261] RLNet 2024 ECCV 2D CNN Voxel 70.88 69.43 78.12 72.81 90.82 78.71 91.67 87.07
@@
[262] InterFusion 2022 IROS CNN+RPN Pillar 67.50 63.21 78.79 69.83 88.11 74.80 87.50 83.47
@@
[250] CenterPoint 2021 CVPR VoxelNet Point 32.74 38.00 65.42 45.42 62.01 48.18 84.98 65.66
@@
[255] PointPillars 2019 CVPR 2D-CNN Pillar 65.55 55.71 72.96 64.74 81.10 67.92 88.96 79.33
@@
The ELMAR algorithm[256] is designed for single-frame operation, pairing LiDAR with 4D radar without any temporal accumulation. Its radar branch incorporates a Dynamic Motion-Aware Encoding module. It learns to classify entire objects as moving or static via a dedicated motion-aware loss, avoiding the unreliability of fixed velocity cutoffs. This produces motion descriptors that are stable enough to be used directly in detection. After each branch produces its initial set of bounding boxes, the Cross-Modal Uncertainty Alignment module attempts to match boxes between modalities, using geometric differences between matched pairs as a fine-grained measure of uncertainty. Considerable disagreements often cause predictions to automatically downweigh during training, directing the model towards cases where both sensors agree. This means the detector only captures the advantages of the radar’s motion sensitivity but also uses the radar’s consistency to overcome the LiDAR’s weaknesses in challenging conditions.
@@
The MoRAL framework [257] was developed with the goal of tackling the distortions that often result from combining multiple 4D radar frames, especially the misalignment of fast objects. Conventionally, points from earlier frames can stretch along the direction of motion, which creates elongated ”tails” that distort shape and affect detection efficiency. MoRAL avoids this by embedding a Motion-aware Radar Encoder (MRE) at the front of the radar branch. This module uses a segmentation-based approach to distinguish moving points from static ones, then shifts only the moving points into alignment with a chosen reference frame using their absolute radial velocity and capture time. While the radar branch undergoes this correction, the LiDAR input is processed independently through a sparse encoder. The two streams meet in the Motion Attention Gated Fusion (MAGF) module, which injects motion cues into the LiDAR feature space and applies a gating mask. That mask strengthens the signal from dynamic foreground regions while naturally dampening static background responses before the network predicts final bounding boxes. Meanwhile, L4DR pipeline [258] addresses a different challenge: the steep quality gap between LiDAR and radar data, and the fact that each sensor degrades differently in poor weather. Its processing begins with Foreground-Aware Denoising, which uses learned segmentation to strip radar scans of spurious returns before any fusion occurs. Cleaned radar data is fused early with raw LiDAR through the Multi-Modal Encoder. In this step, points are grouped into shared pillars, and each pillar allows a two-way exchange of information—radar’s velocity and RCS values inform the LiDAR points in that cell. Meanwhile, LiDAR contributes its high-precision geometry and reflectivity back to radar. The fused data then passes into the Inter-modal and Intra-modal backbone, which maintains three streams: one for each modality and one combined. To prevent all three from carrying redundant patterns, a Multi-Scale Gated Fusion block filters each intra-modal stream based on cues from the inter-modal branch, enabling the network to emphasize whichever sensor is more trustworthy at that moment.
@@
[265] Point-Level 2025 MDPI CNN+ECA CNN+ECA - - - 88.34 78.08 75.38 59.67 54.67 50.03 82.49 61.43 57.38
@@
[267] CSMNET 2024 TGRS DeepLabv3+ C-FPS 90.16 85.18 80.51 82.05 72.64 67.10 48.95 42.15 39.44 82.86 67.86 60.90
@@
[268] LoGoNet 2023 CVPR FPN Voxel-RCNN - - - 91.80 85.06 80.74 70.20 63.72 59.46 84.47 71.70 64.67
@@
[269] VirConv-L 2023 CVPR VirConvNet VirConvNet 95.53 91.95 87.07 91.41 85.05 80.22 - - - - - -
@@
[269] VirConv-T 2023 CVPR VirConvNet VirConvNet 96.11 92.65 89.69 92.54 86.25 81.24 73.32 66.93 60.38 90.04 73.90 69.06
@@
[269] VirConv-S 2023 CVPR VirConvNet VirConvNet 95.99 93.52 90.38 92.48 87.20 82.45 - - - - - -
@@
[213] TED-M 2023 AAAI TeSpConv TeSpConv 55.25 79.21 46.52 91.61 85.28 80.68 55.85 49.21 46.52 88.82 74.12 66.84
@@
[270] DFIM 2023 Elsevier 2D CNN SA 92.61 88.69 85.77 88.36 81.37 76.71 49.91 41.97 39.27 76.99 59.12 52.97
@@
[271] MVMM 2023 IEEE ResNet-50 3D CNN 92.17 88.70 85.47 87.59 78.87 73.78 47.54 40.49 38.36 77.82 64.81 58.79
@@
[272] SupFusion-v1 2023 ICCV MVXNet MVXNet - - - 88.20 78.46 74.01 63.11 58.13 53.73 74.79 57.60 53.82
@@
[272] SupFusion-v2 2023 ICCV PV-RCNN PV-RCNN - - - 92.02 85.20 82.94 74.29 66.67 61.38 91.00 76.13 71.66
@@
[273] VPFNet 2022 IEEE 2D Backbone Voxel-RCNN 93.02 91.86 86.94 91.02 83.21 78.20 71.47 63.59 56.85 87.77 71.99 66.14
@@
[274] SFD 2022 CVPR CPConvs Voxel-RCNN 95.64 91.85 86.83 91.73 84.76 77.92 72.94 66.69 61.59 93.39 72.95 67.26
@@
[247] Focals Conv 2022 CVPR 2D CNN 3D CNN 95.45 91.51 91.21 90.55 82.28 77.59 - - - - - -
@@
[275] VFF 2022 CVPR ResNet-50 PV-RCNN 95.43 91.40 90.66 89.58 81.97 79.17 - - - - - -
@@
[275] VFF 2022 CVPR ResNet-50 Voxel-RCNN 95.65 91.75 91.39 89.50 82.09 79.29 73.26 65.11 60.03 89.40 73.12 69.86
@@
[276] Graph-Vol 2022 ECCV DLA-34 RPN 95.69 90.10 86.85 91.89 83.27 77.78 - - - - - -
@@
[280] EPNet++ 2022 IEEE PointRCNN PointRCNN - - - 91.37 81.96 76.71 52.79 44.38 41.29 76.15 59.71 53.67
@@
[281] CLOCs-v1 2020 IROS 2D CNN 2D CNN 91.16 88.23 82.63 86.38 78.45 72.45 62.54 56.76 52.26 - - -
@@
[281] CLOCs-v2 2020 IROS 2D CNN 2D CNN 92.60 88.99 81.74 87.50 76.68 71.20 60.33 54.17 46.42 - - -
@@
[281] CLOCs-PVCas 2020 IROS 2D CNN 2D CNN 93.05 89.80 86.57 88.94 80.67 77.15 - - - - - -
@@
[282] EPNet 2020 ECCV 2D CNN SA + FP 94.22 88.47 83.69 89.81 79.28 74.59 52.79 44.38 41.29 76.15 59.71 53.67
@@
[283] PointPainting-v1 2020 CVPR DeepLabv3+ PointPillars - - - 90.01 87.65 85.65 77.25 72.41 68.53 81.72 68.76 63.99
@@
[283] PointPainting-v2 2020 CVPR DeepLabv3+ VoxelNet - - - 90.05 87.51 86.66 73.16 65.05 57.33 87.46 68.08 65.59
@@
[283] PointPainting-v3 2020 CVPR DeepLabv3+ PointRCNN 92.13 87.39 82.72 90.19 87.64 86.71 72.65 66.06 61.24 86.33 73.69 70.17
@@
[285] PointFusion 2018 CVPR ResNet-50 PointNet - - - 77.92 63.00 53.27 33.36 28.04 23.38 49.34 29.42 26.98
@@
[286] F-PointNet -v1 2018 CVPR FPN-Based PointNet 87.28 77.09 67.90 80.62 64.70 56.07 50.88 41.55 38.04 69.36 53.50 52.88
@@
[286] F-PointNet -v2 2018 CVPR FPN-Based PointNet++ 88.70 84.00 75.33 81.20 70.39 62.19 51.21 44.89 40.23 71.96 56.77 50.39
@@
From Table 21, we provide a detailed discussion of four selected models: VirConv [269], obtained the highest results in $AP_{3D}$ CarCar evaluations; ViKIENet [263], which achieved the highest results in $AP_{BEV}$ CarCar evaluations; PointPainting [283], got the highest results in $AP_{3D}$ PedestrainPedestrain evaluations; and SupFusion [272], which attained the highest results in $AP_{3D}$ CyclistCyclist evaluations on the KITTI dataset. VirConv [269] is a multimodal 3D object detection framework that uses virtual points to strengthen LiDAR–RGB fusion while tackling two persistent issues in depth-completion-based augmentation: excessive point density and noise from inaccurate depth estimates. Its core component, the VirConv operator, combines Stochastic Voxel Discard (StVD) to filter out redundant near-range voxels while preserving geometry-rich distant points, and Noise-Resistant Submanifold Convolution (NRConv) to encode features jointly in 3D space and their 2D image projections, enabling effective suppression of boundary noise without eroding useful edge detail. From this foundation, three detector configurations were developed. VirConv-L applies an early-fusion scheme aimed at minimizing processing time. VirConv-T follows a late-fusion route and adds multi-transformation refinement to raise detection precision. VirConv-S extends the approach into a semi-supervised framework, using high-confidence pseudo-labels.

### v2

@@ 5 Object Detection in Autonomous Vehicles
[bai2022transfusion] TransFusion 2022 CVPR 3D CNN Point 70.2 65.5 86.2 56.7 66.3 58.8 86.1 68.3 44.2
@@
[fazlali2022versatile] VMVF 2022 CVPR 3D CNN Range 66.3 58.3 84.4 51.5 58.3 53.9 85.3 63.0 32.5
@@
[yin2021center] CenterPoint 2021 CVPR VoxelNet Point 65.5 58.0 84.6 51.0 60.2 53.2 83.4 53.7 28.7
@@
[pan20213d] Pointformer 2021 CVPR Pointformer Point 63.8 53.8 82.3 45.1 48.3 45.4 85.1 55.7 25.8
@@
[chen2020every] CVCNET 2020 NeurIPS 3D CNN Voxel 66.6 58.2 82.6 49.5 59.4 51.1 83.0 61.8 38.8
@@
[yang20203dssd] 3DSSD 2020 CVPR SA Point 56.4 42.6 81.2 47.2 61.4 30.5 70.2 36.0 8.6
@@
[yin2020lidar] 3DVID 2020 CVPR PMPNet Pillar 53.1 45.4 79.7 33.6 47.1 43.1 76.5 40.7 19.9
@@
[zhu2019class] CBGS 2019 CVPR Sparse CNN Voxel 63.3 52.8 81.1 48.5 54.9 42.9 80.1 51.5 22.3
@@
[lang2019pointpillars] PointPillars 2019 CVPR 2D CNN Pillar 45.3 30.5 68.4 23.0 28.2 23.4 59.7 27.4 1.1
@@
Table 19 presents a comparative analysis of 3D object detection methods using the LiDAR-based NuScenes dataset. From Table 19, we provide a detailed discussion of three selected models: PointPillars [lang2019pointpillars], recognized as one of the pioneer model in this field; Voxel Mamba [zhang2024voxel], which achieved the highest results in terms of NDS (NuScenes Detection Score), mAP, and trailer detection on the NuScenes dataset; and LION [liu2024lion], obtained the highest results for car, truck, bus, pedestrian, and cyclist detection on NuScenes dataset. PointPillars [lang2019pointpillars] adopts a pillar-based representation to enable a purely 2D convolutional pipeline for real-time 3D object detection. The approach first discretizes the point cloud into vertical columns (pillars) with infinite extent along the z-axis, then uses a simplified PointNet to extract point-wise features within each pillar. These are aggregated and scattered into a dense pseudo-image in the bird’s-eye-view plane. A 2D CNN backbone, composed of a top-down path and a multi-scale upsampling path, processes this pseudo-image to extract high-level spatial features. Finally, an SSD-style detection head performs classification and oriented 3D bounding box regression directly in BEV space, allowing end-to-end training without any 3D convolutions.
@@
Voxel Mamba [zhang2024voxel] introduces a group-free state-space backbone for 3D object detection, exploiting the linear-time sequence modeling of State Space Models (SSMs) to bypass the computational bottlenecks of Transformer-based voxel encoders. Instead of partitioning the voxel grid into smaller local groups, it serializes all non-empty voxels into a single sequence using a Hilbert Input Layer, which maps 3D voxel coordinates into a 1D order while preserving spatial locality. The backbone employs a Dual-scale SSM Block (DSB), wherein a forward branch processes the high-resolution sequence. In contrast, a backward branch operates on a down-sampled BEV representation, extending the receptive field without quadratic complexity. An Implicit Window Partition (IWP) injects relative position embeddings for intra- and inter-window offsets, enabling the model to recover local context cues without explicit grouping. LION [liu2024lion] replaces transformer-based voxel backbones with linear recurrent neural operators to model long-range dependencies across large voxel regions while avoiding quadratic complexity. Point clouds are voxelized and organized into fixed-size 3D windows, within which LION Blocks process voxel features hierarchically. Each block contains a pair of linear group RNN operators, applied along orthogonal spatial axes, to propagate information efficiently across extended spatial extents. Before this recurrent mixing, a 3D spatial descriptor based on submanifold convolutions enriches local geometric context to offset the positional information loss during voxel flattening. Within its hierarchy, the model reduces voxel resolution at selected stages, representing features at multiple scale, and giving them higher-level layers access.
@@
Table 20 presents a comparative analysis of 3D object detection methods using LiDAR-based data from the VOD dataset. From Table 20, we provide a detailed discussion of three selected models: ELMAR [peng2025elmar] and MoRAL [peng2025moral], which achieved the highest results in terms of $AP\%$ for entire area and driving corridor on VOD dataset; and L4DAR [huang2025l4dr], recognized as one of the notable model in the context of 3D object detection.
@@
[peng2025elmar] ELMAR 2025 ArXiv DS+SA Point 76.41 69.34 78.91 74.89 91.74 81.35 93.01 88.70
@@
[peng2025moral] MoRAL 2025 ArXiv Sparse CNN Pillar 71.23 69.67 79.01 73.30 90.91 78.90 96.25 88.68
@@
[peng2025mutualforce] MutualForce 2025 ICASSP CNN+RPN Pillar 71.67 66.26 77.35 71.76 92.31 76.79 89.97 86.36
@@
[huang2025l4dr] L4DR 2024 AAAI PointNet++ Pillar 69.10 66.20 82.80 72.70 90.80 76.10 95.50 87.47
@@
[deng2024robust] CM-FA 2024 ICRA SA Point 71.39 68.54 76.60 72.18 90.91 80.78 87.80 86.50
@@
[xu2024rlnet] RLNet 2024 ECCV 2D CNN Voxel 70.88 69.43 78.12 72.81 90.82 78.71 91.67 87.07
@@
[wang2022interfusion] InterFusion 2022 IROS CNN+RPN Pillar 67.50 63.21 78.79 69.83 88.11 74.80 87.50 83.47
@@
[yin2021center] CenterPoint 2021 CVPR VoxelNet Point 32.74 38.00 65.42 45.42 62.01 48.18 84.98 65.66
@@
[lang2019pointpillars] PointPillars 2019 CVPR 2D-CNN Pillar 65.55 55.71 72.96 64.74 81.10 67.92 88.96 79.33
@@
The ELMAR algorithm[peng2025elmar] is designed for single-frame operation, pairing LiDAR with 4D radar without any temporal accumulation. Its radar branch incorporates a Dynamic Motion-Aware Encoding module. It learns to classify entire objects as moving or static via a dedicated motion-aware loss, avoiding the unreliability of fixed velocity cutoffs. This produces motion descriptors that are stable enough to be used directly in detection. After each branch produces its initial set of bounding boxes, the Cross-Modal Uncertainty Alignment module attempts to match boxes between modalities, using geometric differences between matched pairs as a fine-grained measure of uncertainty. Considerable disagreements often cause predictions to automatically downweigh during training, directing the model towards cases where both sensors agree. This means the detector only captures the advantages of the radar’s motion sensitivity but also uses the radar’s consistency to overcome the LiDAR’s weaknesses in challenging conditions.
@@
The MoRAL framework [peng2025moral] was developed with the goal of tackling the distortions that often result from combining multiple 4D radar frames, especially the misalignment of fast objects. Conventionally, points from earlier frames can stretch along the direction of motion, which creates elongated ”tails” that distort shape and affect detection efficiency. MoRAL avoids this by embedding a Motion-aware Radar Encoder (MRE) at the front of the radar branch. This module uses a segmentation-based approach to distinguish moving points from static ones, then shifts only the moving points into alignment with a chosen reference frame using their absolute radial velocity and capture time. While the radar branch undergoes this correction, the LiDAR input is processed independently through a sparse encoder. The two streams meet in the Motion Attention Gated Fusion (MAGF) module, which injects motion cues into the LiDAR feature space and applies a gating mask. That mask strengthens the signal from dynamic foreground regions while naturally dampening static background responses before the network predicts final bounding boxes. Meanwhile, L4DR pipeline [huang2025l4dr] addresses a different challenge: the steep quality gap between LiDAR and radar data, and the fact that each sensor degrades differently in poor weather. Its processing begins with Foreground-Aware Denoising, which uses learned segmentation to strip radar scans of spurious returns before any fusion occurs. Cleaned radar data is fused early with raw LiDAR through the Multi-Modal Encoder. In this step, points are grouped into shared pillars, and each pillar allows a two-way exchange of information—radar’s velocity and RCS values inform the LiDAR points in that cell. Meanwhile, LiDAR contributes its high-precision geometry and reflectivity back to radar. The fused data then passes into the Inter-modal and Intra-modal backbone, which maintains three streams: one for each modality and one combined. To prevent all three from carrying redundant patterns, a Multi-Scale Gated Fusion block filters each intra-modal stream based on cues from the inter-modal branch, enabling the network to emphasize whichever sensor is more trustworthy at that moment.
@@
[shen2025point] Point-Level 2025 MDPI CNN+ECA CNN+ECA - - - 88.34 78.08 75.38 59.67 54.67 50.03 82.49 61.43 57.38
@@
[uzair2024channel] CSMNET 2024 TGRS DeepLabv3+ C-FPS 90.16 85.18 80.51 82.05 72.64 67.10 48.95 42.15 39.44 82.86 67.86 60.90
@@
[li2023logonet] LoGoNet 2023 CVPR FPN Voxel-RCNN - - - 91.80 85.06 80.74 70.20 63.72 59.46 84.47 71.70 64.67
@@
[wu2023virtual] VirConv-L 2023 CVPR VirConvNet VirConvNet 95.53 91.95 87.07 91.41 85.05 80.22 - - - - - -
@@
[wu2023virtual] VirConv-T 2023 CVPR VirConvNet VirConvNet 96.11 92.65 89.69 92.54 86.25 81.24 73.32 66.93 60.38 90.04 73.90 69.06
@@
[wu2023virtual] VirConv-S 2023 CVPR VirConvNet VirConvNet 95.99 93.52 90.38 92.48 87.20 82.45 - - - - - -
@@
[wu2023transformation] TED-M 2023 AAAI TeSpConv TeSpConv 55.25 79.21 46.52 91.61 85.28 80.68 55.85 49.21 46.52 88.82 74.12 66.84
@@
[chen2023lidar] DFIM 2023 Elsevier 2D CNN SA 92.61 88.69 85.77 88.36 81.37 76.71 49.91 41.97 39.27 76.99 59.12 52.97
@@
[li2023mvmm] MVMM 2023 IEEE ResNet-50 3D CNN 92.17 88.70 85.47 87.59 78.87 73.78 47.54 40.49 38.36 77.82 64.81 58.79
@@
[qin2023supfusion] SupFusion-v1 2023 ICCV MVXNet MVXNet - - - 88.20 78.46 74.01 63.11 58.13 53.73 74.79 57.60 53.82
@@
[qin2023supfusion] SupFusion-v2 2023 ICCV PV-RCNN PV-RCNN - - - 92.02 85.20 82.94 74.29 66.67 61.38 91.00 76.13 71.66
@@
[zhu2022vpfnet] VPFNet 2022 IEEE 2D Backbone Voxel-RCNN 93.02 91.86 86.94 91.02 83.21 78.20 71.47 63.59 56.85 87.77 71.99 66.14
@@
[wu2022sparse] SFD 2022 CVPR CPConvs Voxel-RCNN 95.64 91.85 86.83 91.73 84.76 77.92 72.94 66.69 61.59 93.39 72.95 67.26
@@
[chen2022focal] Focals Conv 2022 CVPR 2D CNN 3D CNN 95.45 91.51 91.21 90.55 82.28 77.59 - - - - - -
@@
[li2022voxel] VFF 2022 CVPR ResNet-50 PV-RCNN 95.43 91.40 90.66 89.58 81.97 79.17 - - - - - -
@@
[li2022voxel] VFF 2022 CVPR ResNet-50 Voxel-RCNN 95.65 91.75 91.39 89.50 82.09 79.29 73.26 65.11 60.03 89.40 73.12 69.86
@@
[yang2022graph] Graph-Vol 2022 ECCV DLA-34 RPN 95.69 90.10 86.85 91.89 83.27 77.78 - - - - - -
@@
[liu2022epnet++] EPNet++ 2022 IEEE PointRCNN PointRCNN - - - 91.37 81.96 76.71 52.79 44.38 41.29 76.15 59.71 53.67
@@
[pang2020clocs] CLOCs-v1 2020 IROS 2D CNN 2D CNN 91.16 88.23 82.63 86.38 78.45 72.45 62.54 56.76 52.26 - - -
@@
[pang2020clocs] CLOCs-v2 2020 IROS 2D CNN 2D CNN 92.60 88.99 81.74 87.50 76.68 71.20 60.33 54.17 46.42 - - -
@@
[pang2020clocs] CLOCs-PVCas 2020 IROS 2D CNN 2D CNN 93.05 89.80 86.57 88.94 80.67 77.15 - - - - - -
@@
[huang2020epnet] EPNet 2020 ECCV 2D CNN SA + FP 94.22 88.47 83.69 89.81 79.28 74.59 52.79 44.38 41.29 76.15 59.71 53.67
@@
[vora2020pointpainting] PointPainting-v1 2020 CVPR DeepLabv3+ PointPillars - - - 90.01 87.65 85.65 77.25 72.41 68.53 81.72 68.76 63.99
@@
[vora2020pointpainting] PointPainting-v2 2020 CVPR DeepLabv3+ VoxelNet - - - 90.05 87.51 86.66 73.16 65.05 57.33 87.46 68.08 65.59
@@
[vora2020pointpainting] PointPainting-v3 2020 CVPR DeepLabv3+ PointRCNN 92.13 87.39 82.72 90.19 87.64 86.71 72.65 66.06 61.24 86.33 73.69 70.17
@@
[xu2018pointfusion] PointFusion 2018 CVPR ResNet-50 PointNet - - - 77.92 63.00 53.27 33.36 28.04 23.38 49.34 29.42 26.98
@@
[qi2018frustum] F-PointNet -v1 2018 CVPR FPN-Based PointNet 87.28 77.09 67.90 80.62 64.70 56.07 50.88 41.55 38.04 69.36 53.50 52.88
@@
[qi2018frustum] F-PointNet -v2 2018 CVPR FPN-Based PointNet++ 88.70 84.00 75.33 81.20 70.39 62.19 51.21 44.89 40.23 71.96 56.77 50.39
@@
From Table 21, we provide a detailed discussion of four selected models: VirConv [wu2023virtual], obtained the highest results in $AP_{3D}$ CarCar evaluations; ViKIENet [yu2025vikienet], which achieved the highest results in $AP_{BEV}$ CarCar evaluations; PointPainting [vora2020pointpainting], got the highest results in $AP_{3D}$ PedestrainPedestrain evaluations; and SupFusion [qin2023supfusion], which attained the highest results in $AP_{3D}$ CyclistCyclist evaluations on the KITTI dataset. VirConv [wu2023virtual] is a multimodal 3D object detection framework that uses virtual points to strengthen LiDAR–RGB fusion while tackling two persistent issues in depth-completion-based augmentation: excessive point density and noise from inaccurate depth estimates. Its core component, the VirConv operator, combines Stochastic Voxel Discard (StVD) to filter out redundant near-range voxels while preserving geometry-rich distant points, and Noise-Resistant Submanifold Convolution (NRConv) to encode features jointly in 3D space and their 2D image projections, enabling effective suppression of boundary noise without eroding useful edge detail. From this foundation, three detector configurations were developed. VirConv-L applies an early-fusion scheme aimed at minimizing processing time. VirConv-T follows a late-fusion route and adds multi-transformation refinement to raise detection precision. VirConv-S extends the approach into a semi-supervised framework, using high-confidence pseudo-labels.